# Invariant derivations and trace bounds

### v3

@@ 13. Components other than the identity
Fix a coset representative $w_{0}$, so that $w=w_{0}t$ with $t$ uniform random from $T$. We thus see in particular that $\operatorname{Ad}(w)|_{\mathfrak{t}}=\operatorname{Ad}(w_{0})|_{\mathfrak{t}}$, so the claimed expectation and variance are indeed constants.
Now, as a function of $t\in T$, $\operatorname{Tr}\operatorname{Ad}(w)-\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})$ is a Laurent polynomial, which we may compute as a sum over the orbits of $w_{0}$ on the roots. A non-singleton orbit contributes 00 to the trace, while the trace on a root space $\mathfrak{t}_{\alpha}$ fixed by $w_{0}$ is a scalar of norm 1 times the monomial corresponding to $\alpha$. The first claim then follows because every monomial that appears is nonconstant, and thus is uniformly distributed in $S^{1}$. For the second claim, we use Parseval’s theorem to compute
$\mathbb{E}\left[\left|\operatorname{Tr}(w)-\operatorname{Tr}(w|_{\mathfrak{t}})\right|^{2}\right]=\sum_{\alpha:w_{0}\alpha=\alpha}\left|\operatorname{Tr}(w)|_{\mathfrak{t}_{\alpha}}\right|^{2}=\sum_{\alpha:w_{0}\alpha=\alpha}1.\qed$
@@
For any coset $w_{0}T$ of $T$ in $N_{G}(T)$, we have the inequalities
$\min_{w\in w_{0}T}\operatorname{Tr}(\operatorname{Ad}(w))\leq\operatorname{Tr}(\operatorname{Ad}(w_{0})|_{\mathfrak{t}})\leq\max_{w\in w_{0}T}\operatorname{Tr}(\operatorname{Ad}(w)),$
with each equality holding if and only if $w_{0}$ fixes no roots.
@@
The inequalities follow from connectivity of $T$ and the mean value theorem. If $w_{0}$ fixes no roots, then the trace is constant on $w_{0}T$, so equality holds; if $w_{0}$ fixes a root, then the real random variable $\operatorname{Tr}\operatorname{Ad}(w_{0}t)$ has nonzero variance and thus attains values strictly below and above its mean. ∎
@@
For any component $C$ of $G$ and $w_{0}\in N_{G}(T)\cap C$,
$\min_{g\in C}\operatorname{Tr}\operatorname{Ad}(g)\leq\operatorname{Tr}(\operatorname{Ad}(w_{0})|_{\mathfrak{t}})\leq\max_{g\in C}\operatorname{Tr}\operatorname{Ad}(g).$
If either equality holds, then $w_{0}$ fixes no roots.
@@
Since $w_{0}T\subset C$, we have
$\min_{g\in C}\operatorname{Tr}\operatorname{Ad}(g)\leq\min_{w\in w_{0}T}\operatorname{Tr}(\operatorname{Ad}(w))\leq\operatorname{Tr}(\operatorname{Ad}(w_{0})|_{\mathfrak{t}})\\ \leq\max_{w\in w_{0}T}\operatorname{Tr}(\operatorname{Ad}(w))\leq\max_{g\in C}\operatorname{Tr}\operatorname{Ad}(g),$
establishing the first claim. For the second claim, note that if the given equality holds, then so does the corresponding equality in the previous corollary, and thus $w_{0}$ fixes no roots as required. ∎
@@
Let $x\in N_{G}(T)$. Then among $w\in xN_{G^{\circ}}(T)$, there is a choice $w_{0}$ that fixes no roots of $G$ with respect to $T$ and satisfies
(15.5) $\inf_{g\in xG^{\circ}}\operatorname{Tr}\operatorname{Ad}(g)=\operatorname{Tr}(w_{0}|_{\mathfrak{t}}).$
If $G^{\circ}$ is simple, then $w_{0}$ can be chosen to additionally satisfy the following equivalent properties:
@@
    $\mathfrak{t}^{w_{0}}=0$, i.e., $1$ is not an eigenvalue for $w$ acting on $\mathfrak{t}$.
@@
Next we note that any of the three equivalent conditions implies that $w_{0}$ fixes no roots of $G$. Indeed, if $w_{0}$ fixes a root $\alpha$, then it fixes the corresponding coroot $\alpha^{\vee}$ in the dual root system, then the corresponding cocharacter $\mathbb{G}_{\mathrm{m}}\to T$, which would imply $\mathfrak{t}^{w_{0}}\neq 0$.
@@
Simple case: Suppose first that the identity component $G^{\circ}$ of $G$ is simple; this is the crux case. If $-1\in xN_{G^{\circ}}(T)/T$, we can take $w_{0}=-1$; it fixes no roots and $\operatorname{Tr}(w_{0}|_{\mathfrak{t}})=-(\operatorname{rank}G)$ is the minimum of $\operatorname{Tr}\operatorname{Ad}(g)$ by Theorem 2.3a. So suppose $-1\notin xN_{G^{\circ}}(T)/T$, and in particular $G^{\circ}$ has type $A_{n}$ for $n\geq 2$, $D_{n}$ for $n\geq 4$, or $E_{6}$. As in the statement of Theorem 2.3, conjugation by $x$ induces a graph automorphism $s$ of the Dynkin diagram, and we refer to Theorem 2.3 for the value of $\underline{m}(s)$.
@@
If $G^{\circ}$ has type $A$, then it is isogenous to $\operatorname{SU}(n)$ for some $n\geq 3$ and $s=1$. Take $w_{0}$ to be any $n$-cycle in the Weyl group $S_{n}$. The action of the Weyl group on the coroot space is equivalent to the subspace of the permutation representation of $S_{n}$ consisting of vectors whose coordinates sum to 1. Any fixed vector in $\mathfrak{t}$ then must have all coordinates equal and therefore must be 0. Since $w_{0}$ acts on the permutation representation with trace zero, it has trace $-1=\underline{m}(s)$ on $\mathfrak{t}$.
@@
If $G^{\circ}$ has type $D_{n}$ for odd $n\geq 5$, then $s$ has order 1 and $\underline{m}(s)=2-n$. We view $G$ as $\operatorname{SO}_{2n}$, which contains a copy of $\operatorname{SO}_{6}\times\operatorname{SO}_{2n-6}$ of type $A_{3}\times D_{n-3}$. Take $w_{0}$ to be a 4-cycle in the Weyl group of $\operatorname{SO}_{6}=\operatorname{SL}_{4}/\mu_{2}$ and $-1$ in the Weyl group of $\operatorname{SO}_{2n-6}$. The first part has trace $-1$ on its 3-dimensional subalgebra of $\mathfrak{t}$ and the second has trace $3-n$ on its $(n-3)$-dimensional subalgebra of $\mathfrak{t}$, for a total trace of $2-n=\underline{m}(s)$. By the same argument as in the previous paragraph, $w_{0}$ does not fix any vector in $\mathfrak{t}$.
@@
If $G^{\circ}$ has type $D_{n}$ for even $n\geq 4$ and $s$ has order 2, then after multiplying $x$ by an element of $G^{\circ}$ we may assume that conjugation by $x$ normalizes the set of simple roots and so is a graph automorphism of the Dynkin diagram order 2. Composing conjugation by $x$ with the Weyl group element that is $-1$ on the torus gives an element $w_{0}$ that evidently fixes no roots and has $\operatorname{Tr}(w|_{\mathfrak{t}})=2-n=\underline{m}(s)$.
@@
If $G^{\circ}$ has type $D_{4}$ and $s$ has order 3, then the subgroup of $N_{G}(T)/T$ generated by $xT$ and the Weyl group has order 576 and contains two conjugacy classes of elements of order 3 that have exactly 8 elements. One of these two classes is in $xN_{G^{\circ}(T)}/T$. Any representative of that class provides an element $w_{0}$ that acts on the root space $\mathfrak{t}$ with trace $-2=\underline{m}(s)$ and does not have 1 as an eigenvalue.
@@
If $G^{\circ}$ has type $E_{6}$, take $w_{0}$ to be an element of order 3. There are three conjugacy classes of such elements in the Weyl group, see for example [Car72, Table 9]. We take $w_{0}$ to be in the unique conjugacy class of size 80. In the action on the root space $\mathfrak{t}$, $w_{0}$ has trace $-3=\underline{m}(s)$ and does not have 1 as an eigenvalue.
@@
For the general case, conjugation by $x$ permutes the simple factors of $G^{\circ}$. Put $G^{\circ}\cong G_{1}\times G_{2}$, where $G_{1}$ is a product of the central torus and the simple factors that are normalized by $x$ and $G_{2}$ is the a product of the simple factors that are not normalized by $x$. Now every $g\in xG^{\circ}$ has $\operatorname{Tr}\operatorname{Ad}(g)$ equal to the trace of the projection of $\operatorname{Ad}(g)$ on $\operatorname{Lie}(G_{1})$ and every $w\in xN_{G^{\circ}}(T)/T$ moves every root in $G_{2}$ and $\operatorname{Tr}(w|_{\operatorname{Lie}(T)})$ equals the trace of the restriction of $w$ to $\operatorname{Lie}(T\cap G_{1})$. So it suffices to apply the lemma for the case $G^{\circ}=G_{1}$ and extend the resulting $w_{0}$ in any way to an element of $xN_{G^{\circ}}(T)/T$. ∎
@@
Take $x\in G$ minimizing $\operatorname{Tr}\operatorname{Ad}(x)$. Then $X$ normalizes some maximal torus $T$ of $G$ [Bou05, §IX.5.3, Th. 1(b)]. Apply Lemma 15.4 to produce a $w_{0}\in xG^{\circ}$ with $\operatorname{Tr}\operatorname{Ad}(x)=\operatorname{Tr}(w_{0}|_{\mathfrak{t}})$, so
$\min_{g\in G}\operatorname{Tr}\operatorname{Ad}(g)\geq\min_{w\in W}\operatorname{Tr}(w|_{\operatorname{Lie}(T)}).$
@@ 16. Minimum values of characters: $\operatorname{SU}(2)$
Perhaps the most natural collection of representations to consider next are the absolutely irreducible representations of $\operatorname{SU}(2)$. Put $\chi_{d}$ for the character of the irreducible (complex) representation $\operatorname{Sym}^{d}(\mathbb{C}^{2})$ with highest weight $d$, i.e., $s^{d}+s^{d-2}+\cdots+s^{-(d-2)}+s^{-d}$. In the notation of Example 5.1 and (9.1), we have $t=\chi_{1}=s+1/s$ and $\chi_{d}\in\mathbb{C}[T]^{W}=\mathbb{C}[t]$. That is, $\chi_{d}$ can be written as a polynomial in $t$, which we denote by $\chi_{d}^{t}$, where we have already written the first few cases in (9.1). It is easy to see that
@@
We can provide a uniform bound. Recall that $\chi_{d}$ denotes the function on $\operatorname{SU}(2)$.
@@ 17. What kind of numbers are these?
For $G$ a compact real Lie group, what sort of real number might $\min_{g\in G}\operatorname{Re}\chi(g)$ be? When $G$ is simple and $\chi=\operatorname{Tr}\operatorname{Ad}$ or the trace of the highest short root representation, we showed that it is an integer (Theorem 2.3 and Proposition 2.5). For the same $G$ and general $\chi$, since elements of finite order are dense in $G$, it follows that the minimum is a limit of algebraic integers. Examples 8.1, 8.2, and 16.2 exhibit cases where $\min_{g\in G}\operatorname{Re}\chi(g)$ is an algebraic number, but not an algebraic integer. We have the following general result.
@@
Let $\rho$ be a representation of a compact real Lie group $G$. For each connected component $C$ of $G$, the extreme values of $\operatorname{Re}\operatorname{Tr}\rho$ on $C$ are algebraic numbers.
@@
Let $T=T_{c}$ and $c$ be as in Lemma 13.1. Since $C$ is a union of conjugates of $cT$, the extreme values of $\operatorname{Re}\operatorname{Tr}\rho$ are obtained on $cT$. Replacing $G$ by the subgroup generated by $c$ and $T$, we may assume that $G$ is abelian and $G^{\circ}$ is a torus, i.e., a product $U(1)_{\mathbb{R}}^{r}$ of $r$ copies of $U(1)_{\mathbb{R}}$ for some $r$.
We claim that $G$, viewed as an algebraic group, is defined over the algebraic closure $R$ of $\mathbb{Q}$ in $\mathbb{R}$. For $m$ the order of $c$ in $G$, we have an exact sequence
@@
where $U(1)_{\mathbb{R}}^{r}$ is identified with $G^{\circ}$. The elements of the kernel $K$ have coordinates in $U(1)_{\mathbb{R}}$ that are elements of order dividing $m$, which all belong to $U(1)_{R}$, verifying the claim. This argument also shows that the connected component $C$ of $G$ is defined over $R$.
@@
Next we note that $\rho$ is also defined over $R$. Since $G$ is abelian, $\rho$ is a sum of one-dimensional representations, i.e., it amounts to a homomorphism $G\to U(1)_{\mathbb{R}}^{n}$ for some $n$. Its restriction to $T$ is a homomorphism $U(1)_{\mathbb{R}}^{r}\to U(1)_{\mathbb{R}}^{n}$, such that composing this map with projection on any of the $n$ coordinates gives a map of the form $(z_{1},z_{2},\ldots,z_{r})\mapsto\prod_{j=1}^{n}z_{j}^{e_{j}}$ for integers $e_{j}$ [Bou05, §IX.7]. If we re-write this product, expanding $z_{j}=x_{j}+iy_{j}$ for $x_{j},y_{j}$ real, we find that the real and imaginary parts of $\prod z_{j}^{e_{j}}$ are each given by a polynomial in the $x_{j}$ and $y_{j}$ with integer coefficients. That is, $\rho|_{T}$ arises by base change from a polynomial function $U(1)_{R}^{r}\to U(1)_{R}^{n}$. The projection of $\rho(c)$ in each of the $n$ coordinates is an $m$-th root of unity, where $m$ is the order of $c$ in $G$, so $\rho(c)$ is an element of $U(1)_{R}^{n}$.
@@
The trace provides a morphism of schemes from $U(1)_{R}^{n}$ to the Weil restriction $\operatorname{Res}_{\mathbb{Q}^{\mathrm{a}}/R}(\mathbb{A}^{1}_{R})$, whose $R$-points equal the algebraic closure $\mathbb{Q}^{\mathrm{a}}$ of $\mathbb{Q}$. This algebraic group is naturally identified with pairs of elements $(z,\bar{z})$ for $z\in\mathbb{Q}^{\mathrm{a}}$, so it is clear that there is furthermore a morphism of schemes from it to $\mathbb{A}^{1}_{R}$. The composition of $\rho$ with these maps provides a morphism of schemes $G\to\mathbb{A}^{1}_{R}$ whose base change to $\mathbb{R}$ is $\operatorname{Re}\operatorname{Tr}\rho$. We restrict this map to the connected component $C$ of $G$. Its graph is a closed subscheme of $C\times\mathbb{A}^{1}_{R}$. The projection of the graph on the second component — i.e., the image of $\operatorname{Re}\operatorname{Tr}\rho$ — is a semi-algebraic subset of $\mathbb{A}^{1}_{R}$ by the Tarski-Seidenberg Projection Theorem [Sch24, Th. 1.5.8].
@@
Let us now consider the image of $\operatorname{Re}\operatorname{Tr}\rho$ over $\mathbb{R}$. Because $C\times\mathbb{R}$ is compact, the image is a finite union of closed intervals in $\mathbb{R}$. In particular the boundary of the image is a finite set of points including the extreme values. The Tarski Transfer Principle implies then that the boundary of the image over the subfield $R$ is a semi-algebraic set over $R$ [Sch24, Cor. 1.6.17]. Therefore the extreme values are algebraic over $R$, proving the claim. ∎
@@
The second and third paragraphs of the proof, showing that $G$ and $\rho$ are defined over $\mathbb{R}$, could be replaced by various other arguments. For example, one could use the classification of compact Lie groups to express $G$ as obtained by Galois descent from a reductive complex algebraic group whose identity component is split. Repeating this same descent for the quadratic extension $\mathbb{Q}^{\mathrm{a}}$ of $R$ shows that $G$ is defined over $R$, and a similar argument would show the same for $\rho$.

### v4

@@ 13. Components other than the identity
First suppose that $w=1$. Write the Lie algebra $\mathfrak{g}$ of $G$ as a sum of weight spaces with respect to $T$: $\mathfrak{g}=\mathfrak{t}\oplus\left(\oplus_{\alpha\in\Phi}\mathfrak{g}_{\alpha}\right)$ where $\Phi:=\{\alpha\in T^{*}\mid\mathfrak{g}_{\alpha}\neq 0\}$ is the set of roots of $G$. The restriction of each $\alpha$ to a rank 1 subtorus $S^{1}$ of $T$ is uniformly distributed, so has mean 0. Since $T$ is isomorphic to a product of copies of $S^{1}$, it follows that $\operatorname{Tr}(\operatorname{Ad}_{G}(t)|_{\mathfrak{g}_{\alpha}})$ has mean 0 and therefore that $\operatorname{Tr}(\operatorname{Ad}_{G}(t))$ has mean $\operatorname{rank}(G)$ as claimed. Regarding the variance, $\operatorname{Tr}(\operatorname{Ad}_{G}(t))$ has Fourier expansion $\operatorname{rank}(G)+\sum_{\alpha\in\Phi}\alpha(t)$, so Parseval’s Theorem gives
$\mathbb{E}\left[\left|\operatorname{Tr}(\operatorname{Ad}_{G}(t))-\operatorname{rank}(G)\right|^{2}\right]=\sum_{\alpha\in\Phi}\left|1\right|^{2}=\left|\Phi\right|$
as claimed.
Recall the subquotient $H$ defined just before Lemma 13.2; it depends on the coset $wT$. We next treat the case where $w$ centralizes $H$. By Lemma 13.4, we have
(15.2) $\operatorname{Tr}(\operatorname{Ad}_{G}(wt))-\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})=\operatorname{Tr}\operatorname{Ad}_{H}(t)-\operatorname{rank}(H)$
for $t\in T$. Because $T$ surjects on to a maximal torus of $H$, we may apply the $w=1$ case to the right side of (15.2) to see that the left side has mean 0. Neither $\operatorname{Tr}(\operatorname{Ad}_{G}(w)|_{\mathfrak{t}})$ nor $\operatorname{rank}(H)$ depend on $t$, so the variance of $\operatorname{Tr}(\operatorname{Ad}_{G}(wt))$ is the variance of either side of (15.2), which is in turn the variance of $\operatorname{Tr}(\operatorname{Ad}_{H}(t))$, which by the $w=1$ case equals the number of roots of $H$, i.e., the number of roots fixed by $w$, as claimed.
Finally we consider the general case. By Lemma 13.2, there is some $w^{\prime\prime}$ in $wT$ such that $w^{\prime\prime}$ centralizes $H$. Since $w^{\prime\prime}$ belongs to the same coset $wT$, we have $\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})=\operatorname{Tr}(\operatorname{Ad}(w^{\prime\prime})|_{\mathfrak{t}})$ and $w$ and $w^{\prime\prime}$ fix the same roots of $G$. ∎
@@
For any coset $wT$ of $T$ in $N_{G}(T)$, we have the inequalities
$\min_{w^{\prime}\in wT}\operatorname{Tr}(\operatorname{Ad}(w^{\prime}))\leq\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})\leq\max_{w^{\prime}\in wT}\operatorname{Tr}(\operatorname{Ad}(w^{\prime})),$
with each equality holding if and only if $w$ fixes no roots.
@@
The inequalities follow from connectivity of $T$ and the mean value theorem. If $w$ fixes no roots, then the trace is constant on $wT$, so equality holds; if $w$ fixes a root, then the real random variable $\operatorname{Tr}\operatorname{Ad}(wt)$ has nonzero variance and thus attains values strictly below and above its mean. ∎
@@
For any component $C$ of $G$ and $w\in N_{G}(T)\cap C$,
$\min_{g\in C}\operatorname{Tr}\operatorname{Ad}(g)\leq\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})\leq\max_{g\in C}\operatorname{Tr}\operatorname{Ad}(g).$
If either equality holds, then $w$ fixes no roots.
@@
Since $wT\subset C$, we have
$\min_{g\in C}\operatorname{Tr}\operatorname{Ad}(g)\leq\min_{w^{\prime}\in wT}\operatorname{Tr}(\operatorname{Ad}(w^{\prime}))\leq\operatorname{Tr}(\operatorname{Ad}(w)|_{\mathfrak{t}})\\ \leq\max_{w^{\prime}\in wT}\operatorname{Tr}(\operatorname{Ad}(w^{\prime}))\leq\max_{g\in C}\operatorname{Tr}\operatorname{Ad}(g),$
establishing the first claim. For the second claim, note that if the given equality holds, then so does the corresponding equality in the previous corollary, and thus $w$ fixes no roots as required. ∎
@@
Let $x\in N_{G}(T)$. There exists $w\in xN_{G^{\circ}}(T)$ that fixes no roots of $G$ with respect to $T$ and satisfies
(15.6) $\inf_{g\in xG^{\circ}}\operatorname{Tr}\operatorname{Ad}(g)=\operatorname{Tr}(w|_{\mathfrak{t}}).$
If $G^{\circ}$ is simple, then $w$ can be chosen to additionally satisfy the following equivalent properties:
@@
    $\mathfrak{t}^{w}=0$, i.e., $1$ is not an eigenvalue for $w$ acting on $\mathfrak{t}$.
@@
Next we prove that any of the three equivalent conditions implies that $w$ fixes no roots of $G$. Indeed, if $w$ fixes a root $\alpha$, then it fixes the corresponding coroot $\alpha^{\vee}$ in the dual root system, then the corresponding cocharacter $\mathbb{G}_{\mathrm{m}}\to T$, which would imply $\mathfrak{t}^{w}\neq 0$.
@@
Simple case: Suppose first that the identity component $G^{\circ}$ of $G$ is simple; this is the crux case. If $-1\in xN_{G^{\circ}}(T)/T$, we can take $w=-1$; it fixes no roots and $\operatorname{Tr}(w|_{\mathfrak{t}})=-(\operatorname{rank}G)$ is the minimum of $\operatorname{Tr}\operatorname{Ad}(g)$ by Theorem 2.3a. So suppose $-1\notin xN_{G^{\circ}}(T)/T$, and in particular $G^{\circ}$ has type $A_{n}$ for $n\geq 2$, $D_{n}$ for $n\geq 4$, or $E_{6}$. As in the statement of Theorem 2.3, conjugation by $x$ induces a graph automorphism $s$ of the Dynkin diagram, and we refer to Theorem 2.3 for the value of $\underline{m}(s)$.
@@
If $G^{\circ}$ has type $A$, then it is isogenous to $\operatorname{SU}(n)$ for some $n\geq 3$ and $s=1$. Take $w$ to be any $n$-cycle in the Weyl group $S_{n}$. The action of the Weyl group on the coroot space is equivalent to the subspace of the permutation representation of $S_{n}$ consisting of vectors whose coordinates sum to 1. Any fixed vector in $\mathfrak{t}$ then must have all coordinates equal and therefore must be 0. Since $w$ acts on the permutation representation with trace zero, it has trace $-1=\underline{m}(s)$ on $\mathfrak{t}$.
@@
If $G^{\circ}$ has type $D_{n}$ for odd $n\geq 5$, then $s$ has order 1 and $\underline{m}(s)=2-n$. We view $G$ as $\operatorname{SO}_{2n}$, which contains a copy of $\operatorname{SO}_{6}\times\operatorname{SO}_{2n-6}$ of type $A_{3}\times D_{n-3}$. Take $w$ to be a 4-cycle in the Weyl group of $\operatorname{SO}_{6}=\operatorname{SL}_{4}/\mu_{2}$ and $-1$ in the Weyl group of $\operatorname{SO}_{2n-6}$. The first part has trace $-1$ on its 3-dimensional subalgebra of $\mathfrak{t}$ and the second has trace $3-n$ on its $(n-3)$-dimensional subalgebra of $\mathfrak{t}$, for a total trace of $2-n=\underline{m}(s)$. By the same argument as in the previous paragraph, $w$ does not fix any vector in $\mathfrak{t}$.
@@
If $G^{\circ}$ has type $D_{n}$ for even $n\geq 4$ and $s$ has order 2, then after multiplying $x$ by an element of $G^{\circ}$ we may assume that conjugation by $x$ normalizes the set of simple roots and so is a graph automorphism of the Dynkin diagram order 2. Composing conjugation by $x$ with the Weyl group element that is $-1$ on the torus gives an element $w$ that evidently fixes no roots and has $\operatorname{Tr}(w|_{\mathfrak{t}})=2-n=\underline{m}(s)$.
@@
If $G^{\circ}$ has type $D_{4}$ and $s$ has order 3, then the subgroup of $N_{G}(T)/T$ generated by $xT$ and the Weyl group has order 576 and contains two conjugacy classes of elements of order 3 that have exactly 8 elements. One of these two classes is in $xN_{G^{\circ}(T)}/T$. Any representative of that class provides an element $w$ that acts on the root space $\mathfrak{t}$ with trace $-2=\underline{m}(s)$ and does not have 1 as an eigenvalue.
@@
If $G^{\circ}$ has type $E_{6}$, take $w$ to be an element of order 3. There are three conjugacy classes of such elements in the Weyl group, see for example [Car72, Table 9]. We take $w$ to be in the unique conjugacy class of size 80. In the action on the root space $\mathfrak{t}$, $w$ has trace $-3=\underline{m}(s)$ and does not have 1 as an eigenvalue.
@@
For the general case, conjugation by $x$ permutes the simple factors of $G^{\circ}$. Put $G^{\circ}\cong G_{1}\times G_{2}$, where $G_{1}$ is a product of the central torus and the simple factors that are normalized by $x$ and $G_{2}$ is the a product of the simple factors that are not normalized by $x$. Now every $g\in xG^{\circ}$ has $\operatorname{Tr}\operatorname{Ad}(g)$ equal to the trace of the projection of $\operatorname{Ad}(g)$ on $\operatorname{Lie}(G_{1})$ and every $w\in xN_{G^{\circ}}(T)/T$ moves every root in $G_{2}$ and $\operatorname{Tr}(w|_{\operatorname{Lie}(T)})$ equals the trace of the restriction of $w$ to $\operatorname{Lie}(T\cap G_{1})$. So it suffices to apply the lemma for the case $G^{\circ}=G_{1}$ and extend the resulting $w$ in any way to an element of $xN_{G^{\circ}}(T)/T$. ∎
@@
Take $x\in G$ minimizing $\operatorname{Tr}\operatorname{Ad}(x)$. Then $x$ normalizes some maximal torus $T$ of $G$ [Bou05, §IX.5.3, Th. 1(b)]. Apply Lemma 15.5 to produce a $w\in xG^{\circ}$ with $\operatorname{Tr}\operatorname{Ad}(x)=\operatorname{Tr}(w|_{\mathfrak{t}})$, so
$\min_{g\in G}\operatorname{Tr}\operatorname{Ad}(g)\geq\min_{w\in W}\operatorname{Tr}(w|_{\mathfrak{t}}).$
@@ 16. Minimum values of characters: $\operatorname{SU}(2)$
Perhaps the most natural collection of representations to consider next are the absolutely irreducible representations of $\operatorname{SU}(2)$. Put $\chi_{d}$ for the character of the irreducible (complex) representation $\operatorname{Sym}^{d}(\mathbb{C}^{2})$ with highest weight $d$, i.e., $s^{d}+s^{d-2}+\cdots+s^{-(d-2)}+s^{-d}$. In the notation of Example 5.1 and (9.1), we have $t=\chi_{1}=s+s^{-1}$ and $\chi_{d}\in\mathbb{C}[T]^{W}=\mathbb{C}[t]$. That is, $\chi_{d}$ can be written as a polynomial in $t$, which we denote by $\chi_{d}^{t}$, where we have already written the first few cases in (9.1). It is easy to see that
@@
We can provide a uniform bound.
@@ 17. What kind of numbers are these?
For $G$ a compact Lie group, what sort of real number might $\min_{g\in G}\operatorname{Re}\chi(g)$ be? When $G$ is simple and $\chi=\operatorname{Tr}\operatorname{Ad}$ or the trace of the highest short root representation, we showed that it is an integer (Theorem 2.3 and Proposition 2.5). For the same $G$ and general $\chi$, since elements of finite order are dense in $G$, it follows that the minimum is a limit of algebraic integers. Examples 8.1, 8.2, and 16.2 exhibit cases where $\min_{g\in G}\operatorname{Re}\chi(g)$ is an algebraic number, but not an algebraic integer. We have the following general result.
@@
Let $\rho$ be a representation of a compact Lie group $G$. For each connected component $C$ of $G$, the extreme values of $\operatorname{Re}\operatorname{Tr}\rho$ on $C$ are algebraic numbers.
@@
We first set up some notation. Put $\bar{\mathbb{Q}}$ for the algebraic closure of $\mathbb{Q}$ and $F:=\bar{\mathbb{Q}}\cap\mathbb{R}$, the algebraic closure of $\mathbb{Q}$ in $\mathbb{R}$. The unit circle $U(1)$, which we have denoted elsewhere by $S^{1}$ and by which we mean the group scheme over $\mathbb{R}$ whose $\mathbb{C}$-points are the norm 1 complex numbers, is obtained by base change from a group scheme over $F$ we denote by $U(1)_{F}$, whose $F$-points are the norm 1 elements of $\bar{\mathbb{Q}}$.
Pick an element $c\in C$. By the result of de Siebenthal from [dS56] that we used before in the proof of Theorem 2.3 in §13, there is a torus $T$ in $G$ such that $C$ is a union of $G^{\circ}$-conjugates of $cT$. Therefore, the extreme values of $\operatorname{Re}\operatorname{Tr}\rho$ are obtained on $cT$. Replacing $G$ by the subgroup generated by $c$ and $T$, we may assume that $G$ is abelian and $G^{\circ}$ is a torus, i.e., a product $U(1)^{r}$ of $r$ copies of $U(1)$ for some $r$.
We claim that $G$, viewed as an algebraic group, is defined over the algebraic closure $F$ of $\mathbb{Q}$ in $\mathbb{R}$. For $m$ the order of $c$ in $G$, we have an exact sequence
@@
where $U(1)^{r}$ is identified with $G^{\circ}$. The elements of the kernel $K$ have coordinates in $U(1)$ that are elements of order dividing $m$, and in particular these elements already belong to $U(1)_{F}$, verifying the claim. This argument also shows that the connected component $C$ of $G$ is defined over $F$.
@@
Next we prove that $\rho$ is also defined over $F$. Since $G$ is abelian, $\rho$ is a sum of one-dimensional representations, i.e., it amounts to a homomorphism $G\to U(1)^{n}$ for some $n$. Its restriction to $T$ is a homomorphism $U(1)^{r}\to U(1)^{n}$, such that composing this map with projection on any of the $n$ coordinates gives a map of the form $(z_{1},z_{2},\ldots,z_{r})\mapsto\prod_{j=1}^{n}z_{j}^{e_{j}}$ for integers $e_{j}$ [Bou05, §IX.7]. If we re-write this product, expanding $z_{j}=x_{j}+iy_{j}$ for $x_{j},y_{j}$ real, we find that the real and imaginary parts of $\prod z_{j}^{e_{j}}$ are each given by a polynomial in the $x_{j}$ and $y_{j}$ with integer coefficients. That is, $\rho|_{T}$ arises by base change from a polynomial function $U(1)_{F}^{r}\to U(1)_{F}^{n}$. The projection of $\rho(c)$ in each of the $n$ coordinates is an $m$-th root of unity, so $\rho(c)$ is an element of $U(1)_{F}^{n}$.
@@
The trace provides a morphism of schemes from $U(1)_{F}^{n}$ (the norm 1 elements of $\bar{\mathbb{Q}}$) to the Weil restriction $\operatorname{Res}_{\bar{\mathbb{Q}}/F}(\mathbb{A}^{1}_{F})$, whose $F$-points equal the algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$. (The trace does not respect the group operation so this is a morphism of schemes but not group schemes.) This variety is naturally identified with pairs of elements $(z,\bar{z})$ for $z\in\bar{\mathbb{Q}}$, and the map
$\pi(z,\bar{z}):=\tfrac{1}{2}(z+\bar{z})$
is a morphism of schemes $\pi\colon\operatorname{Res}_{\bar{\mathbb{Q}}/F}(\mathbb{A}^{1}_{F})\to\mathbb{A}^{1}_{F}$. The composition $\pi\operatorname{Tr}\rho$ provides a morphism of schemes $G\to\mathbb{A}^{1}_{F}$ whose base change to $\mathbb{R}$ is the map we have denoted elsewhere by $\operatorname{Re}\operatorname{Tr}\rho$. We restrict this map to the connected component $C$ of $G$. Its graph is a closed subscheme of $C\times\mathbb{A}^{1}_{F}$. The projection of the graph on the second component — i.e., the image of $\operatorname{Re}\operatorname{Tr}\rho$ — is a semi-algebraic subset of $\mathbb{A}^{1}_{F}$ by the Tarski-Seidenberg Projection Theorem [Sch24, Th. 1.5.8].
@@
Let us now consider the image of $\operatorname{Re}\operatorname{Tr}\rho$ over $\mathbb{R}$. Because $C$ is compact, its image is a finite union of closed intervals in $\mathbb{R}$. In particular the boundary of the image is a finite set of points including the extreme values. The Tarski Transfer Principle implies then that the boundary of the image over the subfield $F$ is a semi-algebraic set over $F$ [Sch24, Cor. 1.6.18]. Therefore the extreme values are algebraic over $F$, proving the claim. ∎
@@
The second and third paragraphs of the proof, showing that $G$ and $\rho$ are defined over $F$, could be replaced by various other arguments. For example, one could use the classification of compact Lie groups to express $G$ as obtained by Galois descent from a reductive complex algebraic group whose identity component is split. Repeating this same descent for the quadratic extension $\bar{\mathbb{Q}}$ of $F$ shows that $G$ is defined over $F$, and a similar argument would show the same for $\rho$.